OpenStudy (anonymous):

What is the graph of the system y = –2x + 3 and 2x + 4y = –8?

OpenStudy (anonymous):

y = –2x + 3 slope = -2 y-intercept = (0,3) x-intercept = (3/2,0) With two points you can draw a line. . 2x + 4y = –8 4y = -2x -8 y = -(2/4)x -8/4 y = -(1/2)x -2 slope = -1/2 y-intercept = (0,-2) x-intercept = (-4,0)